OpenStudy (anonymous):

How would you do this one (100,-300) And ( 101,-299 )

OpenStudy (katrinakaif):

Your complete question is?

OpenStudy (akshay_budhkar):

Do you need the distance?

OpenStudy (katrinakaif):

Use the same formulas for distance, slope etc.

OpenStudy (anonymous):

find the equestion in slope-intercept form for the line contaning the two poiints (100,-300) And ( 101,-299) Show Work

OpenStudy (anonymous):

But The Problem is i dont get it

OpenStudy (anonymous):

This is what i got Y=-1X+ ? What Would B =

OpenStudy (akshay_budhkar):

ok first u can find the slope using the formula\[m ={ {y2-y1} \over {x1-x2} }\]

OpenStudy (katrinakaif):

First find the slope of the equation using both points.

OpenStudy (katrinakaif):

WIll you set them up for us?

OpenStudy (akshay_budhkar):

ok you need to subsitute wither of the two points in that equation to get the value of b

OpenStudy (katrinakaif):

Use one of the given points and substitute into the equation (y-y1) = m (x-x1) m = slope and solve

OpenStudy (anonymous):

( -300 - - 299 ) ___________________ ( 100- 101 ) = -1 -1 = 1

OpenStudy (anonymous):

??????????